### v1

@@ Introduction
The relationship between environmental enforcement and violence can operate through opposing channels. On one hand, by increasing the cost of deforestation — through a higher probability of fines and apprehensions — enforcement can reduce incentives to engage in violent competition over land and resource extraction. Moreover, if deforestation is an input for other illegal activities, such as mining or constructing airstrips for drug trafficking, enforcement raises the costs of these activities as well. On the other hand, suppressing profitable activities may generate negative income shocks for individuals or groups reliant on them, potentially pushing them toward alternative illicit activities. In such cases, enforcement could inadvertently increase certain forms of violence (MacLeod, 2023; MacLeod and Rivera, 2024).
Understanding which mechanism dominates is crucial for policy design. If environmental enforcement reduces violence, it strengthens the case for integrated strategies that promote both conservation and public security. Conversely, if it exacerbates violence, complementary interventions may be necessary to offset unintended consequences (e.g., as in Soares, 2004; Tuttle, 2019; Deshpande and Mueller-Smith, 2022).
To identify causal effects, we follow the instrumental variable approach proposed by Assunção et al. (2023). Intuitively, we explore exogenous variation in cloud coverage that limits satellite visibility and, in turn, reduces deforestation detection and enforcement intensity in the Brazilian Amazon biome during the stable implementation period of DETER for environmental enforcement (2006–2016). In the first stage, cloud coverage predicts the intensity of enforcement actions, measured by the number of fines issued. In the second stage, the predicted enforcement intensity is used to estimate its effect on municipal homicide rates.
@@
These findings suggest that DETER is cost-effective, even when considering only its crime-reducing benefits. When we scale its estimated benefits by the average enforcement intensity and the Amazon’s population size, we find that DETER prevented 1,477 homicides per year, a 15% reduction from the annual average of 8,790 homicides in the region. Combining this reduction with policy costs and willingness-to-pay estimates from the literature, we find that, even ignoring its environmental objectives, DETER’s law enforcement component alone yields a benefit–cost ratio of at least 3.7. These results underscore the substantial deterrent effect of environmental enforcement in regions characterized by weak institutions and contested land use.
Our results also align with evidence that policies expanding state presence in areas of weak institutional capacity can have large effects on violence. For instance, the retaking of slums previously dominated by drug gangs in Rio de Janeiro reduced homicides by 30% (Ferraz et al., 2023) in neighboring areas, while the creation of militarized police squads in Ceará, one of Brazil’s poorest states, reduced homicides by 57% (Mancha et al., 2025).
@@
This study contributes to three strands of the economics literature. On environmental law enforcement, we highlight a co-benefit: in settings with weak institutions and poorly defined property rights, environmental protection and development need not be opposing goals. Previous work has demonstrated the effectiveness of environmental policies in reducing deforestation in the Amazon (Assunção et al., 2020; Assunção et al., 2023; Assunção et al., 2023; Bragança and Dahis, 2022). However, their broader societal implications remain largely unexplored. By analyzing the impact of environmental enforcement on homicide rates, we document a substantial co-benefit of these policies, showing that efforts to curb illegal deforestation also generate meaningful reductions in local violence. This perspective underscores that environmental enforcement yields multidimensional benefits, which should be incorporated into policy evaluation and cost–benefit analysis.
Second, our paper extends a large literature on crime determinants and deterrence policies. Foundational work has emphasized crime as a rational choice problem shaped by deterrence incentives (Becker, 1968; Ehrlich, 1973), while other studies have provided evidence on the effectiveness of policing and enforcement in reducing crime (Levitt, 2002; Di Tella and Schargrodsky, 2004; Draca et al., 2011). Specifically in the literature on violence in the Brazilian Amazon, Fetzer and Marden (2017) shows that the homologation of indigenous territories reduces violence, while other papers focus on the effects of particular illegal markets, such as the mahogany trade (Chimeli and Soares, 2017), illegal gold mining (Pereira and Pucci, 2022), or drug trafficking (Pereira et al., 2024). In contrast, our analysis evaluates a general environmental enforcement policy implemented throughout the entire Amazon biome without the explicit objective of reducing violence.
Third, this paper contributes to the broader literature on deforestation. This literature has examined determinants of deforestation and evaluated policies to curb it, including carbon taxes (Souza-Rodrigues, 2019; Araujo et al., 2025), trade restrictions (Dominguez-Iino, 2021; Hsiao, 2021; Farrokhi et al., 2025), and credit policies (Assunção et al., 2020). Our study shifts the focus from environmental outcomes to social ones. This perspective broadens the scope of conservation benefits beyond carbon emissions, underscoring that in a weak institutional environment the social value of conservation is larger than previously recognized.
@@
This section provides background on the three core elements of the analysis: violence, deforestation, and environmental policies.111 A detailed discussion of each of these elements is presented in Appendix A. These dimensions are interconnected in the Amazon context, where illicit markets thrive under institutional fragility and limited state capacity.
@@
A central innovation of the PPCDAm was the creation of the Real-Time Detection of Deforestation System (DETER), a satellite-based system developed by the Brazilian Institute for Space Research (INPE) to provide real-time alerts of deforestation activity. By processing high-frequency satellite imagery, DETER identifies recent forest disturbances and produces georeferenced alerts that guide inspection teams in the field (INPE, 2018). Because enforcement is more effective when infractions are caught in the act, timely alerts significantly increase the likelihood of sanctioning violators — raising the perceived probability of detection and thereby strengthening deterrence (Ferreira, 2023). Although cloud cover and limited local capacity still pose challenges, the system has proven effective in reducing deforestation (Assunção et al., 2023).
@@
where $w_{j}$ is the baseline return to activity $j\in\{\text{work},\text{def},\text{crime}\}$; $c_{def},c_{crime}$ are punishments if caught; $\theta,\gamma\in[0,1]$ are the probabilities of detection for deforestation and crime; $\varepsilon^{j}$ are idiosyncratic shocks; $r_{work}\coloneqq w_{work}$, $r_{def}\coloneqq(w_{def}-\theta c_{def})$ and $r_{crime}\coloneqq(w_{crime}-\gamma c_{crime})$.
@@
Measuring crime is inherently challenging due to data limitations, particularly in remote and underserved regions such as the Brazilian Amazon.222 For discussions on the difficulties of crime measurement in the Amazon, see Brazilian Forum on Public Safety (2021). In this context, the homicide rate per 100,000 inhabitants has become a widely accepted measure for violence and criminal activity in empirical research (Chimeli and Soares, 2017; Pereira and Pucci, 2022; de Oliveira and Miranda, 2024). Homicides are more likely to be reported than other crimes and are generally recorded consistently across time and space. The homicide rate of municipality $i$ in year $t$ is calculated as the number of homicides per 100,000 inhabitants.
@@
Homicide data in the Amazon face some challenges. Poor classification of causes of death — especially under “ill-defined” or “undetermined intent” categories — can compromise data quality and contribute to the underreporting of violent deaths (Ipea, 2024). To address this issue, we follow the suggestions by Ipea (2025), adopting an inclusive coding strategy and considering three categories for cause-of-death (ICD-10): assaults by various means (X85–Y09), events of undetermined intent (Y10–Y34), and intentional self-harm (X60–X84). We also show results for alternative categories in Section 7. To compute the homicide rate, we also need municipal population data. We use annual population estimates provided by Brazilian Institute of Geography and Statistics (IBGE). For municipality-years with missing population data, we apply spline interpolation.
@@ 4.4.1 Satellite Controls
To account for potential measurement limitations in satellite-based monitoring, we include two municipality-level indicators of obstructions to PRODES imagery, provided annually by INPE. The first measures the proportion of municipal area covered by clouds during each monitoring period. The second captures the share classified as non-observable due to atmospheric interference, such as cloud shadows or smoke from forest fires. Although the identification strategy relies on variation in DETER visibility, controlling for PRODES obstructions helps address concerns that persistent limitations in satellite-based monitoring — across both systems — may reflect broader structural constraints in state capacity or environmental governance. These constraints could influence the allocation of enforcement resources or correlate with violence-related dynamics. Including these controls helps ensure that the variation in DETER cloud coverage exploited for identification is not confounded by latent monitoring gaps that shape institutional responses or local conflict environments.
@@ 4.4.3 Socioeconomic Controls
Commodity price index. Following Assunção et al. (2015), we construct an output price series designed to capture exogenous variation in demand for agricultural commodities relevant to the region. Building on Assunção et al. (2023), we construct a weighted real price index for six major commodities — beef cattle, soybeans, cassava, rice, corn, and sugarcane.333 Together, these land uses account for nearly 85% of the agricultural land in sample municipalities during the study period. For beef cattle, weights reflect the 2004–2005 average ratio of cattle heads to municipal area in each municipality, using data from the Municipal Livestock Survey (PPM/IBGE). For crops, weights are based on the 2004–2005 average ratio of farmland to municipal area for each commodity $c$ in municipality $i$, using data from the Municipal Crop Production Survey (PAM/IBGE).
@@
$Homicide\_Rate_{i,t}=\delta Enforcement_{i,t-1}+\sum_{k}\theta_{k}\mathbf{X}_{i,t}+\psi_{i}+\lambda_{t}+\varepsilon_{i,t}$ (1)
where $Homicide\_Rate_{i,t}$ is the homicide per 100,000 inhabitants of municipality $i$ in year $t$; $Enforcement_{i,t-1}$ is the the total number of deforestation-related fines issued in municipality $i$ in year t1t-1 and is instrumented by $CloudCover_{i,t-1}$; $\mathbf{X}_{i,t}$ is a vector of $k$ municipality-level controls that includes satellite visibility, weather patterns and socioeconomic factors; $\psi_{i}$ is the municipality fixed effect; $\lambda_{t}$ is the year fixed effect; and $\varepsilon_{i,t}$ is the idiosyncratic error. Standard errors are clustered at the municipality level.
@@
To estimate Equation 1, we instrument the enforcement variable using cloud coverage detected by the DETER system. The first stage of the 2SLS approach captures how adverse monitoring conditions affect the issuance of deforestation-related fines. The first-stage specification is given by:
$Enforcement_{i,t-1}=\beta CloudCover_{i,t-1}+\sum_{k}\gamma_{k}\mathbf{X}_{i,t}+\alpha_{i}+\phi_{t}+\epsilon_{i,t}$ (2)
where $Enforcement_{i,t-1}$ is proxied by the total number of deforestation-related fines issued in municipality $i$ in year t1t-1; $CloudCover_{i,t-1}$ is the average cloud coverage in municipality $i$ in year $t$ measured by DETER; $\mathbf{X}_{i,t}$ is a vector of $k$ municipality-level controls that includes satellite visibility, weather patterns and socioeconomic factors; $\alpha_{i}$ is the municipality fixed effect; $\phi_{t}$ is the year fixed effect; and $\epsilon_{i,t}$ is the idiosyncratic error. Standard errors are clustered at the municipality level.
@@ Results
Section 6.1 presents our main estimates of the causal effect of environmental enforcement on the homicide rate. Moreover, Section 6.2 provides a cost-benefit analysis of the DETER program based on the effect found in Section 6.1. Section 6.3 presents the results of the first-stage regression (Equation (2)) while Section 6.4 discusses the mechanisms that may explain these effects.
@@
To value the estimated reduction, we use the willingness-to-pay estimates by Domínguez and Scartascini (2024). They estimate a willingness to pay of \$152 per person per year for a 20% decrease in homicides. Scaling this estimate to our result implies aggregate benefits of roughly \$2.3 billions annually.
@@
Panel B of Table 1 reports the estimated first-stage coefficients (Equation (2)). Cloud coverage observed by DETER is strongly associated with the number of fines issued. In the main specification (Column (3)), the estimated coefficient is –8.88 and statistically significant at the 1% level, indicating that a one-percentage-point increase in annual cloud cover reduces enforcement intensity by nearly nine fines. The first-stage F-statistic of 16.40 exceeds the conventional threshold of 10 suggested by Stock et al. (2002), alleviating concerns about weak instruments and reinforcing the relevance of cloud coverage as a source of exogenous variation in enforcement intensity.
@@
Third, enforcement actions may disrupt the operation and profitability of illegal economic activities such as land grabbing, unauthorized logging, and unregulated mining. By increasing both the expected costs and logistical challenges of operating outside the law, environmental fines can displace or deter organized groups whose activities often rely on coercion and violence. Chimeli and Soares (2017) document the role of violence in sustaining the illegal mahogany trade in the region, while Soares et al. (2021) and Pereira and Pucci (2022) emphasize the close connection between environmental crimes and local criminal networks. In this view, enforcement shifts the economic incentives of actors who depend on force and impunity to maintain territorial control.
@@
This section evaluates the robustness of the main results through sensitivity analyses and different outcome definitions. In Section 7.1, we examine the validity of the instrument following the sensitivity analysis proposed by Cinelli and Hazlett (2025). In Section 7.2, as a robustness check, we re-estimate the model using an alternative measure of violence restricted to ICD-10 categories X85–Y09, which capture assaults committed through various means, and Y10–Y34, which capture events of undetermined intent. Lastly, in Section 7.3, as an extension, we estimate a set of distributional regressions to assess whether the effects of environmental enforcement extend beyond average homicide rates. Specifically, we examine its impact on two binary outcomes: (i) the incidence of any violence (i.e., a dummy equal to one if the homicide rate is greater than zero), and (ii) the presence of endemic violence (i.e., a dummy equal to one if the homicide rate exceeds 10 homicides per 100,000 inhabitants).
In Appendix C, we present an additional robustness check that controls for other conservation policies coexisting with the DETER program.
@@
In the main specification, the dependent variable includes all deaths classified under ICD-10 codes X85–Y09 (assaults), Y10–Y34 (events of undetermined intent), and X60–X84 (intentional self-harm). This broader definition follows the approach adopted by Ipea (2024, 2025) to better capture lethal violence in contexts where underreporting and misclassifications are prevalent.
A potential concern is that our inclusive classification may inadvertently be capturing an effect on actual intentional self-harm. If this were the case, it would bias our results against our findings. The most plausible channel linking environmental enforcement to intentional self-harm would operate through income: stricter enforcement reduces deforestation, which may lower earnings in affected areas, potentially increasing intentional self-harm. This mechanism would predict that enforcement raises intentional self-harm, not lowers them.
To further assess the robustness of the results to alternative outcome definitions, Table 2 compares the 2SLS estimates using a more restrictive classification. The outcome variable in Column (1) considers only deaths explicitly coded as assaults (ICD-10 X85–Y09) as used by Dix-Carneiro et al. (2018), while the outcome variable in Column (2) broadens the definition by also incorporating events of undetermined intent (ICD-10 Y10–Y34). Lastly, Column (3) corresponds to the main specification with the full set of ICD-10 codes. All regressions include municipality and year fixed effects as well as the full set of controls.
@@
(1) (2) (3)
Lagged Enforcement -0.501 -0.550 -0.728∗∗
(0.292) (0.309) (0.351)
Average homicide rate across municipalities 22.35 24.30 28.16
@@
Observations 5,210 5,210 5,210
Municipalities 521 521 521
Notes: 2SLS coefficients are estimated based on Equation (1) from Section 5. Column (1) includes only assaults by various means (ICD-10 X85–Y09), Column (2) adds events of undetermined intent (ICD-10 Y10–Y34) and Column (3) adds intentional self-harm (ICD-10 X60–X84). “Homicide Rate” is the number of homicides per 100,000 inhabitants. “Lagged Enforcement” refers to the total number of fines issued and serves as a proxy for law enforcement effectiveness. The set of control variables contains PRODES cloud coverage and non-observable (satellite); precipitation and temperature (weather); and commodity index, GDP, population density and Ideb scores (socioeconomic). The dataset is a municipality-by-year panel covering the period 2006-2016. The sample includes all Amazon biome municipalities that exhibited variation in forest cover during the sample period and for which deforestation data are available. Standard errors are clustered at the municipality level and reported in parentheses. Significance: $^{*}p<0.1$, $^{**}p<0.05$, $^{***}p<0.01$.
The results remain consistent across all definitions. In Column (1), the estimated effect of enforcement is –0.501 and statistically significant at the 10% level. This implies that the issuance of one additional deforestation-related fine is associated with a reduction of approximately 0.5 homicides per 100,000 inhabitants, representing a 2.2% decrease relative to the average homicide rate across municipalities. In Column (2), the estimated coefficient becomes –0.550, also statistically significant at the 10% level. This effect implies a 2.3% reduction relative to the average homicide rate across municipalities. Importantly, the relative effects are approximately 2.2%–2.5% in all specifications, and the estimated coefficients are statistically indistinguishable across specifications.

### v2

@@ Introduction
The relationship between environmental enforcement and violence can operate through opposing channels. On the one hand, by increasing the cost of deforestation through a higher probability of fines and apprehensions, enforcement can reduce incentives to engage in violent competition over land and resource extraction. Moreover, if deforestation is an input for other illegal activities, such as mining or constructing airstrips for drug trafficking, enforcement raises the costs of these activities as well. On the other hand, suppressing profitable activities may generate negative income shocks for individuals or groups reliant on them, potentially pushing them toward alternative illicit activities. In such cases, enforcement could inadvertently increase certain forms of violence (MacLeod, 2023; MacLeod and Rivera, 2024).
Understanding which mechanism dominates is crucial for policy design. If environmental enforcement reduces violence, it strengthens the case for integrated strategies that promote both conservation and public security. Conversely, if it exacerbates violence, complementary interventions may be necessary to offset unintended consequences (e.g., Soares, 2004; Tuttle, 2019; Deshpande and Mueller-Smith, 2022).
To identify causal effects, we follow the instrumental variable approach proposed by Assunção et al. (2023). Intuitively, we explore exogenous variation in cloud coverage that limits satellite visibility and, in turn, reduces deforestation detection and enforcement intensity in the Brazilian Amazon biome. We explore this relationship during the stable implementation period of DETER for environmental enforcement (2006–2016) — after which the system was gradually dismantled, particularly under President Bolsonaro’s administration (2019-2022). In the first stage, cloud coverage predicts the intensity of enforcement actions, measured by the number of fines issued. In the second stage, the predicted enforcement intensity is used to estimate its effect on municipal homicide rates.
@@
When we scale these estimated benefits by the average enforcement intensity and the Amazon’s population size, we find that DETER prevented 1,477 homicides per year. This decrease represents a 15% reduction from the annual average of 8,790 homicides in the region. Combining this reduction with policy costs and willingness-to-pay estimates from the literature, we find that, even ignoring its environmental objectives, DETER’s law enforcement component alone yields a benefit–cost ratio of at least 3.7. These results underscore the substantial deterrent effect of environmental enforcement in regions characterized by weak institutions and contested land use.
Our results also align with evidence that policies expanding state presence in areas of weak institutional capacity can have large effects on violence. For instance, the retaking of slums previously dominated by drug gangs in Rio de Janeiro reduced homicides by 30% in neighboring areas (Ferraz et al., 2023), while the creation of militarized police squads in Ceará, one of Brazil’s poorest states, reduced homicides by 57% (Mancha et al., 2025).
@@
This study contributes to three strands of the economics literature. Concerning environmental law enforcement, we highlight a co-benefit: in settings with weak institutions and poorly defined property rights, environmental protection and development need not be opposing goals. Previous work has demonstrated the effectiveness of environmental policies in reducing deforestation in the Amazon (Assunção et al., 2020; Assunção et al., 2023; Assunção et al., 2023; Bragança and Dahis, 2022). However, their broader societal implications remained largely unexplored. By analyzing the impact of environmental enforcement on homicide rates, we document a substantial co-benefit of these policies, showing that efforts to curb illegal deforestation also generate meaningful reductions in local violence. This perspective underscores that environmental enforcement yields multidimensional benefits, which should be incorporated into policy evaluation and cost–benefit analysis.
Second, our paper extends a large literature on crime determinants and deterrence policies. Foundational work has emphasized crime as a rational choice problem shaped by deterrence incentives (Becker, 1968; Ehrlich, 1973), while other studies have provided evidence on the effectiveness of policing and enforcement in reducing crime (Levitt, 2002; Di Tella and Schargrodsky, 2004; Draca et al., 2011). Specifically in the literature on violence in the Brazilian Amazon, Fetzer and Marden (2017) shows that the homologation of indigenous territories reduces violence, while other papers focus on the effects of particular illegal markets, such as the mahogany trade (Chimeli and Soares, 2017), illegal gold mining (Pereira and Pucci, 2022), or drug trafficking (Pereira et al., 2024). In contrast, our analysis evaluates a general anti-deforestation enforcement policy implemented throughout the entire Amazon biome without the explicit objective of reducing violence.
Third, this paper contributes to the broader literature on deforestation. This literature has examined determinants of deforestation and evaluated policies to curb it, including carbon taxes (Souza-Rodrigues, 2019; Araujo et al., 2025), trade restrictions (Dominguez-Iino, 2025; Hsiao, 2025; Farrokhi et al., 2025), and credit policies (Assunção et al., 2020). Our study shifts the focus from environmental outcomes to social ones. This perspective broadens the scope of conservation benefits beyond carbon emissions, underscoring that in a weak institutional environment the social value of conservation is larger than previously recognized. [^1] A complementary literature examines development-oriented policies and finds that, while cash transfer programs may increase deforestation (Rocha and Meyer, 2023), they may reduce deforestation when made conditional on conservation (Wong et al., 2023; Cisneros et al., 2022).
@@
This section provides background on the three core elements of the analysis: violence, deforestation, and environmental policies. [^2] A detailed discussion of each of these elements is presented in Appendix A.
@@
A central innovation of the PPCDAm was the creation of the Real-Time Detection of Deforestation System (DETER), a satellite-based system developed by the Brazilian Institute for Space Research (INPE) to provide real-time alerts of deforestation activity. By processing high-frequency satellite imagery, DETER identifies recent forest disturbances and produces georeferenced alerts that guide inspection teams in the field (INPE, 2018). Because enforcement is more effective when infractions are caught in the act, timely alerts significantly increase the likelihood of sanctioning violators — raising the perceived probability of detection and thereby strengthening deterrence (Ferreira, 2023). Although cloud cover and limited local capacity still pose challenges, the system has proven effective in reducing deforestation (Assunção et al., 2023; Vieira et al., 2023; Picchetti, 2024).
@@
where $w_{j}$ is the baseline return to activity $j\in\{\text{work},\text{def},\text{crime}\}$; $c_{def}$ and $c_{crime}$ are punishments if caught; $\theta,\gamma\in[0,1]$ are the probabilities of detection for deforestation and crime; $\varepsilon^{j}$ are idiosyncratic shocks; let $r_{work}\coloneqq w_{work}$, $r_{def}\coloneqq(w_{def}-\theta c_{def})$ and $r_{crime}\coloneqq(w_{crime}-\gamma c_{crime})$.
@@
Measuring crime is inherently challenging due to data limitations, particularly in remote and underserved regions such as the Brazilian Amazon. [^3] For discussions on the difficulties of crime measurement in the Amazon, see Brazilian Forum on Public Safety (2021). In this context, the homicide rate per 100,000 inhabitants has become a widely accepted measure for violence and criminal activity in empirical research (Chimeli and Soares, 2017; Pereira and Pucci, 2022; de Oliveira and Miranda, 2024). Homicides are more likely to be reported than other crimes and are generally recorded consistently across time and space. The homicide rate of municipality $i$ in year $t$ is calculated as the number of homicides per 100,000 inhabitants.
@@
Homicide data in the Amazon face some challenges. Poor classification of causes of death — especially under “ill-defined” or “undetermined intent” categories — can compromise data quality and contribute to the underreporting of violent deaths (Ipea, 2024). To address this challenge, Ipea (2025) uses the concept of hidden homicides by combining an inclusive coding strategy with a machine learning algorithm. Since data on hidden homicides does not exist at the municipal level for our sample period, we circumvent this issue by adopting an even more inclusive coding strategy and considering three categories for cause-of-death (ICD-10): assaults by various means (X85–Y09), events of undetermined intent (Y10–Y34), and intentional self-harm (X60–X84). We also show results for alternative categories in Section 7. To compute the homicide rate, we also need municipal population data. We use annual population estimates provided by Brazilian Institute of Geography and Statistics (IBGE). For municipality-years with missing population data, we apply spline interpolation.
@@ 4.4.1 Satellite Controls
To account for potential measurement limitations in satellite-based monitoring, we include two municipality-level indicators of obstructions to PRODES imagery, which is the system that provides the official annual measure of deforestation, provided by INPE. The first measures the proportion of municipal area covered by clouds during each monitoring period. The second captures the share classified as non-observable due to atmospheric interference, such as cloud shadows or smoke from forest fires. Although the identification strategy relies on variation in DETER visibility, controlling for PRODES obstructions helps address concerns that persistent limitations in satellite-based monitoring — across both systems — may reflect broader structural constraints in state capacity or environmental governance. These constraints could influence the allocation of enforcement resources or correlate with violence-related dynamics. Including these controls helps ensure that the variation in DETER cloud coverage exploited for identification is not confounded by latent monitoring gaps that shape institutional responses or local conflict environments.
@@ 4.4.3 Socioeconomic Controls
Commodity price index. Following Assunção et al. (2015), we construct an output price series designed to capture exogenous variation in demand for agricultural commodities relevant to the region. Building on Assunção et al. (2023), we construct a weighted real price index for six major commodities — beef cattle, soybeans, cassava, rice, corn, and sugarcane. [^4] Together, these land uses account for nearly 85% of the agricultural land in sample municipalities during the study period. For beef cattle, weights reflect the 2004–2005 average ratio of cattle heads to municipal area in each municipality, using data from the Municipal Livestock Survey (PPM/IBGE). For crops, weights are based on the 2004–2005 average ratio of farmland to municipal area for each commodity $c$ in municipality $i$, using data from the Municipal Crop Production Survey (PAM/IBGE).
@@
$Homicide\_Rate_{i,t}=\delta Enforcement_{i,t-1}+\sum_{k}\theta_{k}\mathbf{X}_{i,t-1}+\psi_{i}+\lambda_{t}+\varepsilon_{i,t}$ (1)
where $Homicide\_Rate_{i,t}$ is the homicide per 100,000 inhabitants of municipality $i$ in year $t$; $Enforcement_{i,t-1}$ is the the total number of deforestation-related fines issued in municipality $i$ in year t1t-1 and is instrumented by $CloudCover_{i,t-1}$; $\mathbf{X}_{i,t-1}$ is a vector of $k$ municipality-level controls that includes satellite visibility, weather patterns and socioeconomic factors; $\psi_{i}$ is the municipality fixed effect; $\lambda_{t}$ is the year fixed effect; and $\varepsilon_{i,t}$ is the idiosyncratic error. Standard errors are clustered at the municipality level.
@@
To estimate Equation (1), we instrument the enforcement variable using cloud coverage detected by the DETER system. The first stage of the 2SLS approach captures how adverse monitoring conditions affect the issuance of deforestation-related fines. The first-stage specification is given by:
$Enforcement_{i,t-1}=\beta CloudCover_{i,t-1}+\sum_{k}\gamma_{k}\mathbf{X}_{i,t-1}+\alpha_{i}+\phi_{t}+\epsilon_{i,t}$ (2)
where $Enforcement_{i,t-1}$ is proxied by the total number of deforestation-related fines issued in municipality $i$ in year t1t-1; $CloudCover_{i,t-1}$ is the average cloud coverage in municipality $i$ in year $t$ measured by DETER; $\mathbf{X}_{i,t-1}$ is a vector of $k$ municipality-level controls that includes satellite visibility, weather patterns and socioeconomic factors; $\alpha_{i}$ is the municipality fixed effect; $\phi_{t}$ is the year fixed effect; and $\epsilon_{i,t}$ is the idiosyncratic error. Standard errors are clustered at the municipality level.
@@ Results
Section 6.1 presents our main estimates of the causal effect of environmental enforcement on the homicide rate. Moreover, Section 6.2 provides a cost-benefit analysis of the DETER program based on the effect found in Section 6.1. Section 6.3 presents the results of the first-stage regression (Equation (2)). Section 6.4 discusses the mechanisms that may explain the effects in the outcome equation. Additionally, Appendix D provides heterogeneity analyses by sex, race and age group, showing that the reduction in homicides is concentrated on black male adults.
@@
To value the estimated reduction, we use the willingness-to-pay estimates by Domínguez and Scartascini (2024). They estimate a willingness to pay of \$152 per person per year for a 20% decrease in homicides. Scaling this estimate to our result implies aggregate benefits of roughly \$2.3 billion annually. [^8] To find this value, multiply \$152 by $\frac{15\%}{20\%}$ and by 20.5 million people.
@@
Panel B of Table 1 reports the estimated first-stage coefficients (Equation (2)). Cloud coverage observed by DETER is strongly associated with the number of fines issued. In the main specification (Column (3)), the estimated coefficient is –8.88 and statistically significant at the 1% level, indicating that a one-percentage-point increase in annual cloud cover reduces enforcement intensity by nearly nine fines. The first-stage F-statistic of 16.40 exceeds the conventional threshold of 10 suggested by Stock et al. (2002), reinforcing the relevance of cloud coverage as a source of exogenous variation in enforcement intensity.
@@
Third, enforcement actions may disrupt the operation and profitability of illegal economic activities such as land grabbing, unauthorized logging, and unregulated mining. By increasing both the expected costs and logistical challenges of operating outside the law, environmental fines can displace or deter organized groups whose activities often rely on coercion and violence. Chimeli and Soares (2017) document the role of violence in sustaining the illegal mahogany trade in the region, while Soares et al. (2021) and Pereira and Pucci (2022) emphasize the close connection between environmental crimes and local criminal networks. In this view, enforcement shifts the economic incentives of actors who depend on force and impunity to maintain territorial control. [^9] Regulatory changes in mining or timber extraction could offer an additional explanation for the effect we document. To address this concern, we: (1) re-estimate our main specification including as a control the interaction between annual gold prices from the World Bank and a dummy for municipal mining activity (as in Pereira and Pucci, 2022) constructed using MapBiomas data. The results remain virtually unchanged (point estimate = –0.72, standard error = 0.35 in the specification of Table 1, Column (3), with the gold–mining interaction included); and (2) include as a control the interaction between annual tropical hardwood log prices from the World Bank and a dummy for mahogany presence from Soares et al. (2021). Again, the results are virtually unchanged (point estimate = –0.79, standard error = 0.37 in the specification of Table 1, Column (3), with the timber–mahogany interaction included).
@@
This section evaluates the robustness of the main results through sensitivity analyses and different outcome definitions. In Section 7.1, we examine the validity of the instrument following the sensitivity analysis proposed by Cinelli and Hazlett (2025). In Section 7.2, as a robustness check, we re-estimate the model using alternative measures of violence. We first restricted the outcome to ICD-10 categories X85–Y09, which capture assaults committed through various means. Next, we extend the definition to include ICD-10 categories Y10–Y34, which capture events of undetermined intent, and, lastly, we incorporate ICD-10 category Y35, which captures legal intervention. Additionally, in Section 7.3, as an extension, we estimate a set of distributional regressions to assess whether the effects of environmental enforcement extend beyond the average homicide rate. Specifically, we examine its impact on two binary outcomes: (i) the incidence of any violence (i.e., a dummy equal to one if the homicide rate is greater than zero), and (ii) the presence of endemic violence (i.e., a dummy equal to one if the homicide rate exceeds 10 homicides per 100,000 inhabitants).
In Appendix E, we present an additional robustness check that controls for other conservation policies coexisting with the DETER program.
@@
In the main specification, the dependent variable includes all deaths classified under ICD-10 codes X85–Y09 (assaults), Y10–Y34 (events of undetermined intent), and X60–X84 (intentional self-harm). This broader definition is inspired by the approach adopted by Ipea (2024, 2025) to better capture lethal violence in contexts where underreporting and misclassifications are prevalent. [^10] Additional robustness checks using other disaggregated homicide categories (firearm assaults, assaults with sharp/blunt objects, and legal intervention) are reported in Appendix C.
A potential concern is that our inclusive classification may inadvertently be capturing an effect on actual intentional self-harm. If this were the case, it would bias our results against our findings. The most plausible channel linking environmental enforcement to intentional self-harm would operate through income: stricter enforcement reduces deforestation, which may lower earnings in affected areas, potentially increasing intentional self-harm. This mechanism would predict that enforcement raises intentional self-harm events, rather than lowering them, as we find in Table 1. Moreover, in Appendix C, we show that the effect of lagged enforcement on intentional self-harm is statistically insignificant and the point estimate is negative.
To further assess the robustness of the results to alternative outcome definitions, Table 2 compares the 2SLS estimates using different homicide classifications. The outcome variable in Column (1) considers only deaths explicitly coded as assaults (ICD-10 X85–Y09) as used by Dix-Carneiro et al. (2018), while the outcome variable in Column (2) broadens the definition by also incorporating events of undetermined intent (ICD-10 Y10–Y34). Column (3) corresponds to the main specification, and Column (4) adds deaths due to legal intervention (ICD-10 Y35). All regressions include municipality and year fixed effects as well as the full set of controls.
@@
(1) (2) (3) (4)
Lagged Enforcement -0.501 -0.550 -0.728∗∗ -0.727∗∗
(0.292) (0.309) (0.351) (0.351)
Average homicide rate across municipalities 22.35 24.30 28.16 28.20
@@
    Y135 (legal intervention)
Observations 5,210 5,210 5,210 5,210
Municipalities 521 521 521 521
Notes: 2SLS coefficients are estimated based on Equation (1) from Section 5. Column (1) includes only assaults by various means (ICD-10 X85–Y09), Column (2) adds events of undetermined intent (ICD-10 Y10–Y34), Column (3) adds intentional self-harm (ICD-10 X60–X84), and Column (4) adds legal intervention (ICD-10 Y-35). “Homicide Rate” is the number of homicides per 100,000 inhabitants. “Lagged Enforcement” refers to the total number of fines issued and serves as a proxy for law enforcement effectiveness. The set of control variables contains PRODES cloud coverage and non-observable (satellite); precipitation and temperature (weather); and commodity index, GDP, population density and Ideb scores (socioeconomic). The dataset is a municipality-by-year panel covering the period 2006-2016. The sample includes all Amazon biome municipalities that exhibited variation in forest cover during the sample period and for which deforestation data are available. Standard errors are clustered at the municipality level and reported in parentheses. Significance: $^{*}p<0.1$, $^{**}p<0.05$, $^{***}p<0.01$.
The results remain consistent across all definitions. In Column (1), the estimated effect of enforcement is –0.501 and statistically significant at the 10% level. This implies that the issuance of one additional deforestation-related fine is associated with a reduction of approximately 0.50 homicides per 100,000 inhabitants, representing a 2.2% decrease relative to the average homicide rate across municipalities. In Column (2), the estimated coefficient becomes –0.55, also statistically significant at the 10% level. This effect implies a 2.3% reduction relative to the average homicide rate across municipalities. In Column (4), the estimated coefficient is -0.73 and remains significant at the 5% level. Relative to the average homicide rate across municipalities, this effect represents a reduction of 2.6%. Importantly, the relative effects are approximately 2.2%–2.6% in all specifications, and the estimated coefficients are statistically indistinguishable across specifications.
@@
Moreover, we separately consider the inclusion of deaths classified under ICD-10 code Y35 (legal intervention) in Appendix C. These cases correspond to killings by state agents in situations where the use of lethal force is deemed lawful, such as armed confrontations with the police. While distinct from conventional homicides, they represent a form of lethal violence that may respond to changes in environmental enforcement through increased interaction between law enforcement and illegal activities. We find no statistically significant effect of environmental enforcement on rates of fatalities in confrontations with state agents. This result is consistent with the institutional context in which IBAMA, rather than police forces, carries out most enforcement activities, and confrontations rarely escalate into lethal violence.